During the 1980s, the average emission of carbon from burning fossil fuels on Earth was 5.439 petagrams (Pg) of carbon per year in the form of CO2.

(a) How many kg of C were placed in the atmosphere each year?
kg

(b) How many kg of CO2 were placed in the atmosphere each year?
kg

(c) A metric ton is 1000 kg. How many metric tons of CO2 were placed in the atmosphere each year?
metric tons / year
If there were 5 billion people on Earth, how many tons of CO2 were produced for each person?
tons per person
Answer #1

(a) mass of carbon added per year = 5.439 pentagram = 5.439×1015 g

= 5.439×1012 × 103 g

= 5.439 ×1012 kg ( Answer)

(b) mass of CO2 added per year = mass of Carbon added per year × ( 44g/mol ÷ 12 g/mol ) = 5.439×1012kg ×(44/12)

= 19.943 ×1012 kg

= 1.994 ×1012 kg (Answer)

(c)

Metric tons of CO2 added per year = 1.994 ×109 × 103 kg

= 1.994 × 109 metric tons / year (answer)

Tons of CO2 produced per person = 1.994 ×109 tons/year ÷ (5×109)

= 0.3989 tons/person (answer)